OpenStudy (anonymous):

(x + 7) (x-1) = (x-1)^2

OpenStudy (anonymous):

I tried foiling both sides idk what I did wrong but do you think you could do it that way pretty please?

OpenStudy (anonymous):

You don't have to FOIL this one, there's a \((x-1)\) on both sides so it's simpler.

OpenStudy (anonymous):

okie let's do it...! as ganeshie8 said we get equation like this after shifting terms (x+7) (x-1)- (x-1)^2=0 now you can see we can take x-1 as a common factor.. after taking x-1 as common factor what do you get ?

OpenStudy (anonymous):

im confused!

OpenStudy (anonymous):

okay just divide both sides of equation (x+7) (x-1)- (x-1)^2=0 what do you get??

OpenStudy (anonymous):

why cant I foil it?

OpenStudy (anonymous):

You can. But it makes the problem needlessly complicated.

OpenStudy (anonymous):

you can foil it but why go to complicated side..?

OpenStudy (anonymous):

okay just divide both sides of equation (x+7) (x-1)- (x-1)^2=0 what do you get??

OpenStudy (anonymous):

\[(x+7)(x-1)-(x-1)^2=0\]Do you see how both terms have \((x-1)\)? You can factor that out. \[(x-1)\bigl(x+7)-(x-1)\bigr) = 0\]

OpenStudy (anonymous):

whait

OpenStudy (anonymous):

wait I wrot

OpenStudy (anonymous):

i wrote the problem down wrong

OpenStudy (anonymous):

it

OpenStudy (anonymous):

(x +7) (x -1) = (x +1)^2

OpenStudy (anonymous):

You already asked that one.

OpenStudy (anonymous):

okie in this you have to expand it...

OpenStudy (anonymous):

i did but i still don't get it

OpenStudy (anonymous):

Did you expand it?

OpenStudy (anonymous):

expand?

OpenStudy (anonymous):

FOIL it!

OpenStudy (anonymous):

oh yah but i did it wrong

OpenStudy (anonymous):

okie tell us what you got.. we will check your answer..

OpenStudy (anonymous):

@lovelike_woee

OpenStudy (anonymous):

6x-8

OpenStudy (anonymous):

you are right what's the matter now? find value of x

OpenStudy (anonymous):

alright thank yuo

OpenStudy (anonymous):

That isn't what I got...

OpenStudy (anonymous):

you are welcome..!

OpenStudy (anonymous):

@SACAPUNTAS check you answer..!

OpenStudy (anonymous):

Ok. \[(x+7)(x-1) - (x+1)^2\]\[(x^2 + 7x - x -7) - (x^2 + x + x + 1)\]\[x^2-x^2+7x-3x+7-1\]\[4x-8\] You messed up a minus sign somewhere I guess.

OpenStudy (anonymous):

And I messed up copying. -7 - 1. But yeah.

OpenStudy (anonymous):

@SACAPUNTAS yep you are right... thanks for correcting bro... you are right...! keep up the good work..! :)

OpenStudy (anonymous):

@lovelike_woee ans must be 4x-8

OpenStudy (anonymous):

yah i figured it out too thanks though!